i have my final coming up and i need help on how to solve this problem. I don't understand it what so ever, please state the steps.
QUESTION: write the equation for the conic section in standard for:
Parabola with vertex at the origin and a directrix y=5. Give the coordinates of the focus.

Answer by richwmiller(17219)

Obviously you have to work backwards from the vertex (0,0) and the directrix y=5
And we have two questions :
What is the equation of the conic section?
What is the focus?
What conic sections have one focus, vertex and directrix?
parabola
the parabola y=x^2 has vertex (0,0)
y=-x^2/20
focus =(0,-5)